OpenStudy (4everaddicted2anime):

Find the surface area of a sphere that has a volume of 288pi cu. in.

OpenStudy (abdullahm):

\(\bf\Large Volume = \frac{4}{3}\pi r^3\) They told us that the volume is \(\bf\Large 288\pi\) We need to plug that in and solve for r!

OpenStudy (4everaddicted2anime):

I need to multiply both sides by 3/4 right?

OpenStudy (abdullahm):

Yes :)

OpenStudy (abdullahm):

And then you need to find the cube root of both sides then. We can already divide pi from both sides

OpenStudy (abdullahm):

\(\bf\Large 288 = \frac{4}{3}r^3\) Multiply both sides by 3/4 and then take the cube root of both sides :)

OpenStudy (4everaddicted2anime):

Okay, I got 216=r^3 r=6

OpenStudy (abdullahm):

That's correct! Now we need to plug in r = 6 into the surface area formula! \(\bf\Large SA = 4\pi r^2\) Plug in r = 6 :)

OpenStudy (4everaddicted2anime):

SA=4pi*6^2 SA=36*4pi SA=144pi right?

OpenStudy (abdullahm):

Correct! :)

OpenStudy (4everaddicted2anime):

Thank you for your help

OpenStudy (abdullahm):

It was my pleasure! :)